OpenStudy (anonymous):

Evaluate the function at the given value of the independent variable and simplify. f(x) = 5x^2 - 5x + 4; f(x - 1)

OpenStudy (anonymous):

f(x) = 5x^2 - 5x + 4; f(x - 1) f(x - 1)= 5(x-1)^2 - 5(x-1) + 4 f(x - 1)= 5(x^2-2x+1) - 5x+5 + 4 f(x - 1)= 5x^2-10x+5 - 5x+5 + 4 f(x - 1)= 5x^2-15x+5 + 14